Selected information regarding a​ company's most recent quarter follows​ (all data in​ thousands).

Cost of goods manufactured ​ $1,590

Gross profit ​ $2,630

Operating expenses ​ $4,050

Ending finished goods inventory ​ $480

Sales revenue ​ $6,080

What was the beginning finished goods​ inventory?
Answer #1

Sales-Cost of goods sold=Gross profit

Cost of goods sold=(6080-2630)=$3450

Cost of goods sold=Cost of goods manufactured+beginning finished goods​ inventory-Ending finished goods inventory

3450=1590+beginning finished goods​ inventory-480

beginning finished goods​ inventory=3450+480-1590

=$2340